OpenStudy (anonymous):

The quotient of a number and 2 is the same as the difference of the number doubled and 3. What is the number?

OpenStudy (johnweldon1993):

x/2 = 2x-3 so first multiply both sides by 2 x = 4x - 6 now subtract 4x from both sides -3x = -6 divide both sides by -3 x = 2 Check 2/2 = (2(2)) - 3 1 = (4-3) 1=1 so 2 is correct :)